Make New Entry,
Make Followup Entry
User name flay
Log entry time 08:02:55 on November 10, 2009
Entry number 298741
keyword=11/10/09 Tuesday Owl Shfit Summary
00:25 Target alarm due to temperature
fluctuation; trying to slowly bring beam back
starting at 60 uA -> 100 uA in 10 uA steps.
01:25 Target is responding well to slow ramp up of beam --
currently at 100 uA.
01:45 Target may have PID issues at 100 uA -- sets off target alarm.
We've decided to run at 90 uA beam for production.
- See HALOG 298690,298691
02:30 Multiple beam trips causing target alarm for past 20 mins.
03:15 Taking 90 uA beam for production runs.
04:00 Production at I ~ 90 uA running smoothly.
06:46 Smooth data taking all night, up until
LHRS coda crash -- no response from TS.
- See HALOG 298727
8:00 RHRS Q2 Liquid He level is low; calling tech-on-call to help fix
the problem; all levels are low -- magnet went down